OpenStudy (anonymous):

Which of the following is the simplified form of the quantity 4 y squared minus 9 over the quantity 2 y squared plus y minus 3

OpenStudy (anonymous):

These are the choices: a. 2y plus 3 over y plus 1, with the restriction y ≠ 3 over 2 and y ≠ –1 b. 2y minus 3 over y minus 1, with the restriction y ≠ –3 over 2 and y ≠ 1 c. 2y plus 3 over y plus 1 , with the restrictions y ≠ 3 over 2 and y ≠ 1 d. 2y minus 3 over y minus 1 , with the restrictions y ≠ –3 over 2 and y ≠ –1 Sooo from what just wrote down it's be b... I think?

OpenStudy (campbell_st):

oops its \[\frac{4y^2 - 9}{2y^2 +y - 3}\] factorise the numerator and denominator \[\frac{(2y -3)(2y + 3)}{(2y + 3 )(y -1)}\] eliminate the common factor \[\frac{2y -3}{y - 1}\] the restictions are y cannot equal 3/2 or 1

OpenStudy (anonymous):

 <--- like that?